# Information-Theoretic Limits of Bistatic Integrated Sensing and Communication

### v3

@@ Abstract
The bistatic integrated sensing and communication (ISAC) system avoids the strong self-interference in a monostatic ISAC system by employing a pair of physically separated sensing transceivers and maintaining the merit of co-designing radar sensing and communication on shared spectrum and hardware. In this paper, we study a bistatic ISAC system with the aim of characterizing its capacity-distortion function, i.e., the maximum achievable communication rate when the sensing accuracy satisfies a given distortion constraint. Based on the proposed information-theoretic model of the bistatic ISAC system, we provide a multi-letter representation of the capacity-distortion function. For ease of calculation, we derive several single-letter lower bounds and upper bounds on the capacity-distortion function. In particular, we provide single-letter representations of the capacity-distortion function for the degraded bistatic ISAC channels. Furthermore, we extend the results to a bistatic ISAC broadcast channel model and obtain the capacity-distortion region with a single-letter characterization in a degraded case. Examples are provided to explicitly illustrate the theoretical results and demonstrate the advantages of ISAC over independent communication and sensing, as well as the benefits of leveraging communication to assist sensing in a bistatic system.
@@ I Introduction
In parallel to the ISAC research centered on wireless communication applications [12, 13, 14, 15, 16, 17, 18, 19, 20], the topic has also been the focus of recent information theoretic research. The authors in [21, 22] studied the monostatic ISAC model from an information theoretic perspective and characterize the optimal tradeoff between the capacity of reliable communication and the distortion of state estimation. In the considered system, an ISAC transmitter aims to transmit messages to a receiver through a memoryless state-dependent channel where the state sequence is i.i.d.; meanwhile, the transmitter also aims to estimate the state sequence of the receiver through the backscattered signal, which is modeled as a causal generalized feedback. Following the seminal work in [22, 21], various information theoretic ISAC works were proposed, which could be generally classified into two frameworks according to the considered channel models:
@@
In this paper, we focus on the bistatic radar system, where the state is i.i.d. over time slots. According to the distribution of transceivers, radar systems can be classified as monostatic, bistatic, and multistatic. In particular, a radar with physically-colocated transmit and receiver arrays is called a monostatic radar, and in many cases the same antenna array is used for both transmitting and receiving. A radar with physically-separated transmitter and receiver arrays is called a bistatic radar. If multiple separated receivers are employed with one transmitter, the radar system is called multistatic [34]. Although the bistatic radar system is generally more complicated to implement than the monostatic system [35], the advantages of bistatic radar in suppressing self-interference, enhancing target detection and localization accuracy, and improving resistance to jamming and stealth capabilities has stimulated continued research. For example, a target that is designed to minimize backscatter by reflecting radar energy in other directions is difficult to be detected by monostatic radar, while it may be easily detected by bistatic radar [36]. Also, interference from the transmit array to the receiver array cannot be ignored in the monostatic radar system, while it is negligible in the bistatic radar system as the transmit and receiver arrays are far apart. A crucial problem associated with bistatic radar systems is the synchronization issue, which requires precise alignment of time and frequency between the physically-separated transmitter and receiver. In such a system, the transmitter and receiver have independent positions and motion states, which may lead to synchronization issues in multiple dimensions such as time, frequency, and phase [37]. Common methods used to address synchronization challenges include global positioning system (GPS) clock synchronization [37, 38], phase-locked loops (PLLs) phase synchronization[39, 40], and adaptive synchronization [41], etc.
Inspired by the appealing benefits of bistatic radar, we study the bistatic ISAC system from an information-theoretic perspective.
@@
As in Fig. 1, the bistatic ISAC system considered consists of an ISAC transmitter (ISAC Tx), a communication receiver (ComRx), and a sensing receiver (SenRx). The ISAC Tx sends a codeword to convey some information to the ComRx who knows the channel states noncausally and perfectly. At the same time, the SenRx at another location receives both the radiated signals of the ISAC Tx and the reflected signals from ComRx to perform the sensing task to estimate the channel states. Given the complexity of information theoretic models, we do not consider synchronization issues in the system, which are assumed to be solved by the aforementioned existing technologies.
It is worth noting that regarding the bistatic ISAC system, the most essential difference between our model and existing model is that in our system considered here, although the SenRx is aware of the codebook, it is unaware of the sent message. As a result, the SenRx cannot directly use the sent messages as in [42] when performing sensing. Therefore, it is expected that in a bistatic ISAC system, the randomness of the communication signal exerts a greater impact on sensing than that in a monostatic ISAC system, implying a different communication-sensing performance tradeoff.
While preparing the manuscript, we notice that the paper [43] studied a ISAC problem similar to ours. The main difference between the paper [43] and ours is the choice of distortion metric. The authors of [43] selected logarithmic loss to measure the quality of a soft estimate, and derived the corresponding capacity-distortion function of the ISAC model. In contrast, we study the capacity-distortion function based on a more general distortion metric, including part of results in [43] as a special case of our results.
Under the information theoretic framework, our main contributions are as follows.
@@
  • We further propose the three single-letter lower bounds based on three decoding-and-estimation (DnE) strategies, i.e., blind estimation, partial-decoding-based estimation, and full-decoding-based estimation. Also, three single-letter upper bounds are proposed by introducing a genie and leveraging Csiszar sum identity and the lossy source coding, respectively. In particular, for the two special cases where the SenRx is a degraded form of the ComRx and the ComRx is a degraded form of the SenRx, we provide the single-letter representations of the capacity-distortion function.
  • The results are extended to a bistatic ISAC broadcast channel model and the capacity-distortion region with a single-letter characterization is obtained when the ComRx is statistically stronger than the SenRx.
@@
The rest of the paper is organized as follows. Section II introduces the system model and defines the capacity-distortion function. Section III gives a multi-letter representation of capacity-distortion function, provides some lower bounds and upper bounds of capacity-distortion function with a single-letter representation, and extends the results to a bistatic ISAC broadcast channel model. Section IV gives the two examples to show explicitly the theoretical results on the capacity-distortion function. Section V concludes this paper.
@@
Upper-case letters represent random variables, and lower-case letters represent their realizations. Calligraphic letters denote sets, e.g., $\mathcal{X}$. $\mathbb{R}$, $\mathbb{R}_{+}$ and $\mathbb{N}_{+}$ represent the sets of real, non-negative real numbers, and non-negative integer, respectively. $X^{n}$ denotes the tuple of random variables $(X_{1},X_{2},\cdots,X_{n})$, $\mathbb{E}[X]$ denotes the expectation for a random variable $X$, and $\oplus$ denotes the modulo-2 sum.
@@ II System Model
In this section, we introduce the bistatic ISAC system model [^1] In this paper, we assume that all the nodes in the system are synchronized., which is modeled as a state-dependent memoryless channel (SDMC) with two receivers. The SDMC model with two receivers consists of four finite sets $\mathcal{X},\mathcal{S},\mathcal{Y},\mathcal{Z}$ and a collection of conditional probability mass functions (pmf) $p\left(y,z|x,s\right)$ over $\mathcal{Y}\times\mathcal{Z}$. The state sequence $S^{n}=(S_{1},\ldots,S_{n})$ is independently and identically distributed (i.i.d.) according to a given state distribution $P_{S}(\cdot)$. We assume $S^{n}$ is perfectly and noncausally available to ComRx but unknown to SenRx.
@@ A $\left(2^{n\mathrm{R}},n\right)$ code for the above SDMC with two receivers consists of
We assume that the message $W$ is uniformly distributed over $\left[1:2^{nR}\right]$. The performance of the decoder is measured by the average probability of error $P_{e}^{(n)}=\mathrm{Pr}\big{\{}\hat{W}\neq W\big{\}}$. The accuracy of the state estimation is measured by the average expected distortion [^2] Since the sensing receiver is unaware of the sent messages, we define the distortion at the sensing receiver by adopting the average expected distortion rather than the maximum distortion relying on the sent messages, where the maximum distortion is defined in Remark 3.
@@ III Main results
By the definition of $C(D)$, we have $C(D_{1})\geq C(D_{2})$ for $D_{1}>D_{2}$. To prove the concavity of $C(D)$, we use the time sharing technique. For each block length $n$, assume without loss of generality that $\alpha n$ is an integer for $\alpha\in[0,1]$ and $R_{1}<C(D_{1}),R_{2}<C(D_{2})$. Assume $\left(2^{\alpha n\mathrm{R_{1}}},\alpha n\right)$ and $\left(2^{(1-\alpha)n\mathrm{R_{2}}},(1-\alpha)n\right)$ are a sequence of codes with rates $R_{1}$ and $R_{2}$, respectively. To transmit a message, for the first $\alpha n$ transmissions, sender transmits a codeword from the first code and for the rest of the transmissions, it transmits a codeword from the second code. Using similar approaches as in [44, Theorem 15.3.2], we can show that the rate-distortion pair $(\alpha{R_{1}}+(1-\alpha){R_{2}},\alpha D_{1}+(1-\alpha)D_{2})$ is achievable. According to the definition of capacity-distortion function $C(D)$, we obtain $C(\alpha D_{1}+(1-\alpha)D_{2})\geq\alpha C(D_{1})+(1-\alpha)C(D_{2})$. ∎
@@ III-A The multi-letter representation of C(D)
where $\hat{S}^{*k}(Z^{k})=\big{(}\hat{S}_{1}^{*}(Z^{k}),\hat{S}_{2}^{*}(Z^{k}),\cdots,\hat{S}_{k}^{*}(Z^{k})\big{)}$ and $\hat{S}_{i}^{*}(Z^{k})=\arg\min_{\hat{s}_{i}}\sum_{s_{i}}p(s_{i}|z^{k})d(s_{i},\hat{s}_{i})$, $i=1,2,\cdots,k$ is the optimal estimator.
@@
where $(a)$ follows from the independence of $Z_{k+1}^{k+l}$ and $S^{k}$ given $Z^{k}$ and $(b)$ follows from the independence of $Y_{k+1}^{k+l}$ and $S^{k}$ given $Y^{k}$. Thus, according to the definition of $C_{k}(D)$, we get that $(k+l)C_{k+l}(D)\geq kC_{k}(D)+lC_{l}(D)$ for any $k,l\in\mathbb{N}_{+}$. Further, using the definition of upper and lower limits we can get the limit of $C_{k}(D)$ exists and $\lim_{k\rightarrow\infty}C_{k}(D)=\sup_{k}C_{k}(D)$.
Proof of achievability: Let $p_{X^{k}}(\cdot)$ be the pmf that achieves $C_{k}(D)$ and also satisfies that
$\mathbb{E}\big{[}d\big{(}S^{k},\hat{S}^{*k}(Z^{k})\big{)}\big{]}\leq kD.$
For code length $n$, set $n=kt+r,r<k$. Fixing a product pmf $\prod_{i=1}^{t}p(x_{(i-1)k+1}^{ik})p(x^{r})$, where $p(x_{(i-1)k+1}^{ik})=p(x^{k}),i=1,2,\dots,t$ and $p(x^{r})$ satisfies $\mathbb{E}\big{[}d(S^{r},\hat{S}^{*r}(Z^{r}))\big{]}\leq rD$. Then, using joint typical decoding at the decoder, we can get the average probability of error $P_{e}^{(n)}$ tends to zero if $R<\frac{1}{k}I\left(X^{k};Y^{k}|S^{k}\right)-\delta(\epsilon)$ hold, where $\delta(\epsilon)\rightarrow 0$ when $n\rightarrow\infty$. Similar to the derivation process of (III-A), we can get the distortion satisfies
@@
$\quad\ kR$ $\leq I\left(W;Y^{k},S^{k}\right)+k\epsilon_{k}$
$=I\left(W;Y^{k}|S^{k}\right)+k\epsilon_{k}$
$\leq I\left(X^{k};Y^{k}|S^{k}\right)+k\epsilon_{k}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}kC_{k}(D)+k\epsilon_{k},$ (8)
@@
From the above derivation, we conclude that in the bistatic ISAC model under consideration, the reason that prevents the single-letter representation of the capacity-distortion function from being obtained is that the optimal estimator is a sequence estimator. In the monostatic ISAC model, the optimal estimator is one-shot since the estimator naturally aware of the sent message $X$ and the Markov chain $(X^{i-1},X_{i+1}^{n},Z^{i-1},Z_{i+1}^{n},\hat{S}_{i})-(X_{i},Z_{i})-S_{i}$ holds, which leads to a single-letter representation of the capacity-distortion function.
@@
where $X^{\prime}$ is stationary and ergodic stochastic process, $Y^{\prime}$ is the output corresponding to $X^{\prime}$ and $S^{\prime}$, and $I(X^{\prime};Y^{\prime}|S^{\prime})=\lim_{k\rightarrow\infty}\frac{1}{k}I(X^{k};Y^{k}|S^{k})$.

### v4

@@ Abstract
Bistatic sensing refers to scenarios where the transmitter (illuminating the target) and the sensing receiver (estimating the target state) are physically separated, in contrast to monostatic sensing, where both functions are co-located. In practical settings, bistatic sensing may be required either due to inherent system constraints or as a means to mitigate the strong self-interference encountered in monostatic configurations. A key practical challenge in bistatic radio-frequency radar systems is the synchronization and calibration of the separate transmitter and sensing receiver. In this paper, we are not concerned with these signal processing aspects and take a complementary information-theoretic perspective on bistatic integrated sensing and communication (ISAC). Namely, we aim to characterize the capacity-distortion function—the fundamental tradeoff between communication capacity and sensing accuracy. We consider a general discrete channel model for a bistatic ISAC system and derive a multi-letter representation of its capacity-distortion function. Then, we establish single-letter upper and lower bounds and provide exact single-letter characterizations for degraded bistatic ISAC channels. Furthermore, we extend our analysis to a bistatic ISAC broadcast channel and derive the capacity-distortion region with a single-letter characterization in the degraded case. Numerical examples illustrate the theoretical results, highlighting the benefits of ISAC over separate communication and sensing, as well as the role of leveraging communication to assist sensing in bistatic systems.
@@ I Introduction
In parallel to the ISAC research centered on wireless communication applications [12, 13, 14, 15, 16, 17, 18, 19, 20], the topic has also been the focus of recent information-theoretic research. The authors in [21, 22] studied the monostatic ISAC model from an information-theoretic perspective and characterize the optimal tradeoff between the capacity of reliable communication and the distortion of state estimation. In the considered system, a transmitter aims to transmit messages to a receiver through a memoryless state-dependent channel where the state sequence is independent and identically distributed (i.i.d.); meanwhile, the transmitter also aims to estimate the state sequence of the receiver through the backscattered signal, which is modeled as a causal generalized feedback. Following the seminal work in [22, 21], various information-theoretic ISAC works were proposed, which could be generally classified into two frameworks according to the considered channel models:
@@
In this paper, we focus on the bistatic sensing system, where the state is i.i.d. over time slots. According to the distribution of transceivers, sensing systems can be classified as monostatic, bistatic, and multistatic. In particular, a sensing system with physically-colocated transmit and receiver antennas is called a monostatic sensing system, and in many cases the same antenna array is used for both transmitting and receiving. A sensing system with physically-separated transmitter and receiver antennas is called a bistatic sensing system. If multiple separated receivers are employed with one transmitter, the sensing system is called multistatic [34]. Although the bistatic sensing system is generally more complicated to implement than the monostatic sensing system [35], the advantages of the bistatic sensing system in suppressing self-interference and enhancing target detection and localization accuracy have stimulated continued research. For example, a target that is designed to minimize backscatter by reflecting radar energy in other directions is difficult to be detected by a monostatic sensing system, while it may be easily detected by a bistatic sensing system [36]. Also, interference from the transmit antenna to the sensing receiver antenna may be significant in the monostatic sensing system, while it is negligible in the bistatic sensing system as the transmitter and sensing receiver are far apart. A crucial problem associated with bistatic sensing systems is the synchronization issue, which requires precise alignment of time and frequency between the physically-separated transmitter and receiver. In such a system, the transmitter and receiver have independent positions and motion states, which may lead to synchronization issues in multiple dimensions such as time, frequency, and phase [37]. Common methods used to address synchronization challenges include global positioning system (GPS) clock synchronization [37, 38], phase-locked loops (PLLs) phase synchronization[39, 40], and adaptive synchronization [41], etc.
Motivated by the appealing benefits of bistatic radar, in this work, we study the bistatic ISAC system from an information-theoretic perspective, along the lines of previous information-theoretic ISAC works reviewed before.
@@
As seen in Fig. 1, the considered bistatic ISAC system consists of a transmitter (ISAC Tx), a communication receiver (ComRx), and a sensing receiver (SenRx). The ISAC Tx sends a codeword to convey some information to the ComRx who knows the channel states perfectly. At the same time, the SenRx at another location receives both the radiated signals of the ISAC Tx and the reflected signals from the ComRx to perform the sensing task to estimate the channel states. For this scenario, we consider a general discrete channel model. As such, synchronization issues are irrelevant to the model we are concerned with. In practice, these issues can be assumed to be solved by the aforementioned existing techniques.
It is worth pointing out that the main conceptual difference between the bistatic ISAC information-theoretic problem considered here and the existing bistatic model considered in [42] is that in our case, the SenRx is aware of the communication codebook, but it is unaware of the sent message (i.e., which codeword is sent). Therefore, it is expected that in a bistatic ISAC system, the randomness of the communication signal exerts a greater impact on sensing than that in a monostatic ISAC system, implying a different communication-sensing performance tradeoff.
Note that the paper [43] followed our preprint and studied a similar ISAC problem. Specifically, the authors of [43] selected logarithmic loss to measure the quality of a soft estimate and derived the corresponding capacity-distortion function of the ISAC model. In contrast, we study the capacity-distortion function based on a more general distortion metric, including part of the results in [43] as a special case of our results.
Under the information-theoretic framework, our main contributions are as follows.
@@
  • We propose a single-letter lower bound based on the superposition coding scheme and partial-decoding-based estimation strategy and a single-letter upper bound by introducing a genie-aided estimator. In particular, for the special case where the channel to SenRx is degraded with respect to the channel to ComRx, we provide the single-letter characterization of the capacity-distortion function by showing that the derived upper and lower bounds coincide.
  • The results are extended to a bistatic ISAC broadcast channel model and the capacity-distortion region with a single-letter characterization is obtained for the degraded case.
@@
The rest of the paper is organized as follows. Section II introduces the system model and defines the capacity-distortion function. Section III gives a multi-letter representation of capacity-distortion function, derives some single-letter lower and upper bounds on the capacity-distortion function and the single-letter characterization of the capacity-distortion function in the degraded case, and extends the results to a bistatic ISAC broadcast channel model. Section IV gives the two examples to show explicitly the theoretical results on the capacity-distortion function. Section V concludes this paper.
@@
Upper-case letters represent random variables, and lower-case letters represent their realizations. Calligraphic letters denote sets, e.g., $\mathcal{X}$. $\mathbb{R}_{+}$ and $\mathbb{N}_{+}$ represent the sets of non-negative real numbers and non-negative integer, respectively. $X^{n}$ denotes the tuple of random variables $(X_{1},X_{2},\cdots,X_{n})$, $\mathbb{E}[X]$ denotes the expectation for a random variable $X$, and $\oplus$ denotes the modulo-2 sum.
@@ II System Model
In this section, we introduce the bistatic ISAC system model, which is modeled as a state-dependent memoryless channel (SDMC) with two receivers. The SDMC model with two receivers consists of four finite sets $\mathcal{X},\mathcal{S},\mathcal{Y},\mathcal{Z}$ and a collection of conditional probability mass functions (pmf) $p\left(y,z|x,s\right)$ over $\mathcal{Y}\times\mathcal{Z}$. The state sequence $S^{n}=(S_{1},\ldots,S_{n})$ is i.i.d. according to a given state distribution $P_{S}(\cdot)$. We assume $S^{n}$ is perfectly and noncausally available to ComRx but unknown to SenRx.
@@ A $\left(2^{n\mathrm{R}},n\right)$ code for the above SDMC with two receivers consists of
We assume that the message $W$ is uniformly distributed over $\left[1:2^{nR}\right]$. The performance of the decoder is measured by the average probability of error $P_{e}^{(n)}=\mathrm{Pr}\big{\{}\hat{W}\neq W\big{\}}$. The accuracy of the state estimation is measured by the average expected distortion [^1] Since the SenRx is unaware of the sent message, we define the distortion at the SenRx by adopting the average expected distortion rather than the maximum distortion relying on the sent message, where the maximum expected distortion is defined in Remark 3.
@@ III Main results
By the definition of $C(D)$, we have $C(D_{1})\geq C(D_{2})$ for $D_{1}>D_{2}$. To prove the concavity of $C(D)$, we use the time sharing technique. For each block length $n$, assume without loss of generality that $\alpha n$ is an integer for $\alpha\in[0,1]$ and $R_{1}<C(D_{1}),R_{2}<C(D_{2})$. Assume $\left(2^{\alpha n\mathrm{R_{1}}},\alpha n\right)$ and $\left(2^{(1-\alpha)n\mathrm{R_{2}}},(1-\alpha)n\right)$ are a sequence of codes with rates $R_{1}$ and $R_{2}$, respectively. To transmit a message, for the first $\alpha n$ transmissions, the sender transmits a codeword from the first code and for the rest of the transmissions, it transmits a codeword from the second code. Using similar approaches as in [44, Theorem 15.3.2], we can show that the rate-distortion pair $(\alpha{R_{1}}+(1-\alpha){R_{2}},\alpha D_{1}+(1-\alpha)D_{2})$ is achievable. According to the definition of capacity-distortion function $C(D)$, we obtain $C(\alpha D_{1}+(1-\alpha)D_{2})\geq\alpha C(D_{1})+(1-\alpha)C(D_{2})$. ∎
@@ III-A The multi-letter representation of C(D)
where $\hat{S}^{*k}(Z^{k})=\big{(}\hat{S}_{1}^{*}(Z^{k}),\hat{S}_{2}^{*}(Z^{k}),\cdots,\hat{S}_{k}^{*}(Z^{k})\big{)}$ and $\hat{S}_{i}^{*}(Z^{k})=\arg\min_{s^{\prime}}\sum_{s_{i}}p(s_{i}|z^{k})d(s_{i},s^{\prime})$, $i=1,2,\cdots,k$ is the optimal estimator.
@@
where $(a)$ follows from the independence of $Z_{k+1}^{k+l}$ and $S^{k}$ given $Z^{k}$ and $(b)$ follows from the independence of $Y_{k+1}^{k+l}$ and $S^{k}$ given $Y^{k}$. Thus, according to the definition of $C_{k}(D)$, we get that $(k+l)C_{k+l}(D)\geq kC_{k}(D)+lC_{l}(D)$ for any $k,l\in\mathbb{N}_{+}$, which means that $kC_{k}(D)$ is superadditive sequence. Further, using the definition of upper and lower limits we get the limit of $C_{k}(D)$ exists and $\lim_{k\rightarrow\infty}C_{k}(D)=\sup C_{k}(D)$.
Proof of achievability: Let $p_{X^{k}}(\cdot)$ be the pmf that achieves $C_{k}(D)$ and also satisfies that $\mathbb{E}\big{[}d\big{(}S^{k},\hat{S}^{*k}(Z^{k})\big{)}\big{]}\leq kD.$ For code length $n$, set $n=kt+r,r<k$. Fixing a product pmf $\prod_{i=1}^{t}p(x_{(i-1)k+1}^{ik})p(x^{r})$, where $p(x_{(i-1)k+1}^{ik})=p(x^{k}),i=1,2,\dots,t$ and $p(x^{r})$ satisfies $\mathbb{E}\big{[}d(S^{r},\hat{S}^{*r}(Z^{r}))\big{]}\leq rD$. Then, using joint typical decoding at the decoder, we get the average probability of error $P_{e}^{(n)}$ tends to zero if $R<\frac{1}{k}I\left(X^{k};Y^{k}|S^{k}\right)-\delta(\epsilon)$ hold, where $\delta(\epsilon)\rightarrow 0$ when $n\rightarrow\infty$. Similar to the derivation process of (III-A), we get the distortion satisfies
@@
$\quad\ kR$ $\leq I\left(W;Y^{k},S^{k}\right)+k\epsilon_{k}=I\left(W;Y^{k}|S^{k}\right)+k\epsilon_{k}$
$\leq I\left(X^{k};Y^{k}|S^{k}\right)+k\epsilon_{k}\stackrel{{\scriptstyle(a)}}{{\leq}}kC_{k}(D)+k\epsilon_{k},$ (8)
@@
From the above derivation, we conclude that in the bistatic ISAC model under consideration, the reason that prevents the single-letter representation of the capacity-distortion function from being obtained is that the optimal estimator is a sequence estimator. In the monostatic ISAC model, the optimal estimator is one-shot, i.e., the estimator estimates the state $s_{i}$ in time slot $i$ only depends on $x_{i}$ and $z_{i}$, since the estimator naturally aware of the sent message $X$ and the Markov chain $(X^{i-1},X_{i+1}^{n},Z^{i-1},Z_{i+1}^{n},\hat{S}_{i})-(X_{i},Z_{i})-S_{i}$ holds, which leads to a single-letter representation of the capacity-distortion function.
@@
where the supremum is takev over all staionary and ergodic processes $X^{\prime}$, $Y^{\prime}$ is the output process obtained by passing $X^{\prime}$ through the bistatic channel model, and $I(X^{\prime};Y^{\prime}|S^{\prime})=\lim_{k\rightarrow\infty}\frac{1}{k}I(X^{k};Y^{k}|S^{k})$.